OpenStudy (nwolf):

ind an equation of the line containing the given pair of points (-7,0) and (0,4) what is the equation of the line y=

OpenStudy (anonymous):

Solve the following for y:\[\frac{y-4}{x-0}=\frac{y-0}{x+7}\]\[y=\frac{4 x}{7}+4 \]

OpenStudy (anonymous):

Open office is not rendering the LaTex expression correctly.

OpenStudy (nwolf):

I had it upside down thanks

OpenStudy (anonymous):

You are welcome.